# Unbounded entropy production and violent fragmentation in long-range interacting super-Tonks-Girardeau systems

### v1

@@ IV Quantum to classical behavior
Before assessing the process of statistical relaxation for general $\alpha$, it is instructive to discuss the Fock space dynamics of the excited metastable sTG phase for a particular case. We select $\alpha$ = 3.0, and corresponding $\epsilon_{\alpha=3.0}$ $\approx 0.1015976$. To understand the dynamical fragmentation during TG to sTG quench, we perform a study with increasing orbital number $M$ and present results for the selected values of $M=5,10,15,20$.
IV.1 Violent fragmentation
Fig. 1 depicts the time evolution of the orbital occupation. The figure shows that initially $(t<0)$ the system is fragmented for all choices of orbital number. The occupation in the natural orbitals changes quite rapidly after the quenching procedure. The occupation in the orbitals with initial significant population decreases, while the one in the initially nearly empty orbitals increases. This trend continues until a saturation point, located approximately at time $t=0.5\bar{t}$ for all values of $M$, where the population of the orbitals coalesces around the value $1/M$. This represents an irregular behavior which we term violent fragmentation, and is the hallmark of a complete saturation of the probed Hilbert space of our quantum simulations. We remark that this type of violent fragmentation is exhibited — practically unchanged — for all values of $M$ up to $M=20$, which represents a computational upper threshold for our simulations given the current state-of-the-art. At the transition time the fully collapsed state always occupies the full available configuration space. The TG to sTG quench can therefore be described as a transition process from quantum to classical behavior of the long-range interacting bosons.
@@ IV.1 Violent fragmentation
IV.2 Fock space delocalization
@@
To further understand the quantum to classical behavior, we visualize the value of the coefficients $C_{\bar{n}}(t)$ after the quench in Fig. 2 We plot $|C_{\bar{n}}(t)|^{2}$ as a function of the index $n$ of the basis states for the final simulation time at $t=3.0\bar{t}$. The number of configurations in the available space for particular choice of $M$ is determined by $N_{\mathrm{conf}}$ by Eq.(4). Thus, $N_{\mathrm{conf}}$ increases exponentially from 126 for $M=5$ to over 40,000 for $M=20$. For each choice of $M$, the initial TG gas undergoes a violent delocalization process in which most of the configurations acquire a comparable weight. This behavior corresponds to the definition of classical gas, which occupies the whole available configuration space.
@@
Figure 2: Configuration space population at time $t=3.0\bar{t}$. s measured from the amplitude squared of the many-body coefficients, for $N=5$ quenched bosons with $\alpha=3.0$, $g_{d}=100.0\to-100.0$ and increasing orbital number $M$. The orbital number is (a) $M=5$, (b) $M=10$, (c) $M=15$, (d) $M=20$.
IV.3 Unbounded Entropy measures
@@
In the dynamical evolution of isolated quantum system, statistical relaxation and thermalization are established by some kind of equilibrium. For interacting many-body systems, statistical relaxation is also related to chaos in the energy spectra [103]. In particular, chaos in quantum interacting systems is defined by the statistics of the eigenstates. The onset of chaos is established by quadratic growth in entropy at very short time, then linear growth before saturation to a maximum entropy states. For time reversal and rotationally invariant systems, the maximum entropy is also determined by the predictions of a Gaussian orthogonal ensemble of random matrices(GOE) [104, 47, 105, 106].
@@
Therefore, to further understand the violent dynamics and its impact on the relaxation process, in Fig 3 we plot the four different entropic quantities introduced in section III ($S_{C}(t)$, $S_{n}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$) as measures of irregular dynamics. We present how the different measures approach the GOE prediction as determined by the number of orbitals used in the computation. For a GOE of random matrices, $I_{C}^{GOE}=D/3$ and $S_{C}^{GOE}=\ln 0.48D$, where $D\times D$ is the dimension of the random matrices [104]. To compare with GOE estimates, we set $D$ as equal to the number of configurations $N_{\mathrm{conf}}$ participating in the dynamics. For occupation entropy $S_{n}$, $D$ is set equal to the number of orbitals $M$ used in the computations. Thus $S_{n}^{GOE}$ = $-\sum_{i=1}^{M}\frac{1}{M}\ln(\frac{1}{M})=\ln(M)$. The GOE estimates are evaluated for different $M$ used in the MCTDH-B treatment.
@@
Fig. 3 shows that statistical relaxation is achieved simultaneously with Fock space delocalization and violent dynamical fragmentation at time $t$=$0.5\bar{t}$. In particular, the results for increasing orbital number further support the thesis that the production of entropy is indeed unbounded. For example, for $M=15$, the number of configurations in the dynamics is $N_{conf}$ = 11’628. The corresponding GOE predictions are $S_{C}^{GOE}$ = 8.627, $S_{n}^{GOE}=2.708$, $I_{C}^{GOE}$= 3’876. We observe that for the prequench state $(t=0)$, all the measures are far enough from 00 indicating that the system initially is fragmented and delocalized. With time, all the entropy measures smoothly increase and then saturate. The saturation values for $M=15$ are $S_{C}^{\mathrm{sat}}\approx 8.94$, $S_{n}^{\mathrm{sat}}\approx 2.71$ and $I_{C}^{\mathrm{sat}}=5^{\prime}820$. These values are indeed close to GOE predictions, however a slight discrepancy exists because the interaction strength needed to exhibit TG limit is very large, albeit not infinite, which is the perquisite for GOE prediction. Repeating the same calculations with $M=20$ yields $N_{\mathrm{conf}}$=42’504, and the corresponding GOE predictions are $S_{C}^{GOE}$ = 9.923, $S_{n}^{GOE}=2.995$, $I_{C}^{GOE}$= 14’168. The corresponding saturation values in MCTDH-B computations are $S_{C}^{\mathrm{sat}}\approx 10.24$, $S_{n}^{\mathrm{sat}}\approx 2.99$ and $I_{C}^{\mathrm{sat}}\approx 21^{\prime}372$. The saturation close to GOE prediction indeed exhibits the hallmark of statistical relaxation. However, the impact of violent fragmentation is visible in the unbound growth of the entropy measures with the number of orbitals.
@@
V Understanding of relaxation process
@@
An intriguing question to understand is the intermediate path from a relaxing state to a fully relaxed one. In this section, we thus analyze the many-body quantum dynamics over the entire range of the exponents $\alpha$ $\epsilon$ $\left[0.5,4\right]$. To understand the relaxation process we thoroughly discuss two typical cases: $\alpha=2.7$ (weakly long-range) and $\alpha=0.7$ (strongly long-range). Besides thoroughly investigating the dynamics of bosons, we also present calculations for spinless fermions to establish the universality of the fragmentation and relaxation processes. For both quantum statistics, we maintain the same protocol in the quench dynamics. We finally summarize the entire results in a parameter-time diagram which depicts the phase boundary between non-relaxed gas and relaxed gas.
V.1 Process of relaxation for bosons
@@
First, we discuss the case for strong long-range interactions with $\alpha=0.7$ and measure all the four entropic quantities $S_{C}(t)$, $S_{n}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$, which are plotted in Fig. 4. We observe a concurring estimate from all the four different measures that the relaxation happens around time $t\approx 0.6\bar{t}$. As expected, the saturation values approach the corresponding GOE predictions. At very small times, we observe a quadratic growth in $S_{c}(t)$ and $S_{C}^{N}(t)$ and both of them exhibit peculiar oscillations before reaching the saturation value. This behavior suggests a highly complex dynamics with several intermediate and rather unstable phases before the entropy eventually settles to the final state. $S_{n}$, instead, exhibits an almost linear increase at short times and smoothly settles to the GOE value. $S_{n}$ is more insensitive to irregularities in the dynamics as it is determined by the fragmentation of the many-body states. In fact, in the previous section we already saw that $S_{n}$ is the entropic quantity that best converges to the GOE result.
@@
The results for weak long-range interactions with $\alpha=2.7$ are presented in Fig. 5. The corresponding relaxation process is less complex than for strong long-range interactions. In the intermediate dynamics, the entropy measures exhibit a single but deeper minimum which is superseded by the eventual saturation to the GOE value. There is no clear signature of quasi equilibrium states for this case, and the time to reach the relaxed state is shorter ($t\approx 0.4\bar{t}$) than for strong long-range interactions.
@@
In Fig. 6, we summarize our results in the form of a parameter-time diagram across the interaction power $\alpha$ and the simulation time. Due to the closest agreement with GOE values for $S_{n}$ and its more regular dynamics that allows to determine the relaxation time more precisely, we will consider it as a good metric and use it as an order parameter in the parameter-time diagram. The diagram shows a clear separation between non-relaxed and relaxed states and we observe an exponential boundary between these two regimes. This suggests very long relaxation times in the limit $\alpha\to 0$, which cannot be reached with our current numerics due to the need of employing increasingly larger system sizes  [^3] We remark that the reduction of entropy exhibited by $\alpha=0.5$ at $t>0.6\bar{t}$ is likely due to self-interference of the long-range tails of the interacting particles. In our numerics, although we employ a harmonic confinement, the system is always defined with periodic boundary conditions. If the system size is not large enough, as it might have happened for the $\alpha=0.5$ results, self-interference patterns across the system boundaries will appear and modify the dynamics.. The parameter-time diagram has more distinct characteristics when $\alpha<1$, i.e. when the range is less than the dimension of the system. For larger values, instead, the exponential boundary gradually loses its structure and the relaxation time is almost independent of the choice of $\alpha$.
Figure 6: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched bosons with $g_{d}=100.0\to-100.0$ and $M=10$. The dashed line indicates the exponential boundary between thermalized and non-thermalized states.
V.2 Process of relaxation for fermions
@@
Figure 9: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched fermions with $g_{d}=100.0\to-100.0$ and $M=10$. The dashed line indicates the exponential boundary between thermalized and non-thermalized states.
VI Control of relaxation
@@
It is an established fact that a (near-)integrable system relaxes in at least two time scales related to the faster and slower relaxation processes [107]. Prethermalization occurs in the intermediate time scale before asymptotic relaxation to an equilibrium state. A necessary condition for thermalization is statistical relaxation in various observables to some kind of equlibrium. However, the observed dynamics in the TG to sTG quench is a highly nonintegrable phenomenon. The dynamics is so complex that different intermediate states may exist, but we can not detect them due to their extremely small lifetimes. It is thus impossible to establish the possibility of any prethermal state in the strongly quenched regime within the current computational limitations.
@@
Nevertheless, another key question that can be addressed is whether it is possible to control the complex and extremely nontrivial pathways observed in the relaxation dynamics during the TG to sTG quench by probing regimes of weaker interactions. In the following, we thus present an extended analysis of the dynamics for quenches with $g_{d}=+100.0$ to $g_{d}=-80.0,-50.0,-30.0$ and $-10.0$. Again, we evaluate the measures of entropy dynamics for $N=5$ bosons for two distinct choices of $\alpha$. We present the results for $\alpha=0.7$ in Fig. 10 and for $\alpha=2.7$ in Fig. 11. For $\alpha=0.7$, we observe that the dynamics is strongly affected by decreasing the interactions to weaker values. The system exhibits more complex pathways and the relaxation process is severely hampered. In the same time scale, when $g_{d}$ is quenched from $+100.0$ to $-100.0$ and the system exhibits perfect relaxation after passing through many complex intermediate states, the $g_{d}$= $+100.0$ to $-10.0$ quench is unable to achieve a completely relaxed state.
It is known that a linear increase in entropy is the precursor of thermalization [109]. In Fig. 10 (b), we indeed observe that $S_{n}(t)$ exhibits a linear increase when the interaction strength is quenched to the same magnitude. Gradually weakening the quench, though, the linear increase starts to disappear and instead a plateau starts to develop. For $g_{d}=+100.0$ to $g_{d}=-10.0$ quench, the growth of $S_{n}(t)$ clearly exhibits three different time scales rather than a linear increase of the entropy. In the first interval (green background), there is an almost steady region. In the second interval (yellow background), $S_{n}(t)$ increases more rapidly and seems to follow logarithmic growth. In the third interval (red background), $S_{n}(t)$ becomes approximately constant and does not quite reach the GOE saturation value. These features indicate the possibility of generating a prethermal state by combining longer-ranged interactions with less violent quenches in generalized long-range systems.
@@
Figure 10: Dynamics for four different entropy measures for $N=5$ quenched, nearly dipolar-interacting bosons with $\alpha=0.7$ and varying quench strength. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$. For all curves $M=10$ orbitals were used. In panel (b), the three different background colors denote the three different regimes observed in the $g_{d}=-100.0\to+10$ quench.
@@
The dynamics shown in Fig. 11 for shorter range interactions and smaller quench is distinctly different. The system remains mostly insensitive to the reduction in the quench strength. Some modulations in the amplitude are observed, but the basic conclusion remains unchanged: the system saturates to the GOE value in a single time scale, irrespective of the post-quench parameters. Fig 11(b) shows that $S_{n}(t)$ maintains its linearity at all times which guarantees thermal-like relaxation for all weaker quench processes. No intermediate plateau is formed. Notably, the saturation value is almost independent on the postquench parameter. We thus conclude that when $\alpha$ is smaller than the dimension of the system, the dynamics can be drastically modified by the strength of the quench protocol, whereas for shorter range interactions the dynamics remains insensitive to it.
@@
VII Conclusions and Outlook
Understanding non-equilibrium dynamics and relaxation of quantum many-body systems is one of the major unsolved problems in quantum mechanics. The scarcity of appropriate theoretical tools to probe the relaxation dynamics makes the scenario even more challenging. When a generic isolated quantum system relaxes, the process of thermalization is not always known. There may be partial relaxation, or different stages of relaxation that can happen on different time scales leading to very complex nonequilibrium dynamics.

### v2

@@ 4 Quantum to classical behavior
Before assessing the process of statistical relaxation for general $\alpha$, it is instructive to discuss the Fock space dynamics of the excited metastable attractive phase for a particular case. We select $\alpha$ = 3.0, and corresponding $\epsilon_{\alpha=3.0}$ $\approx 0.1015976$. To understand the dynamical fragmentation during the quench, we perform a study with increasing orbital number $M$ and present results for the selected values of $M=5,10,15,20$.
4.1 Violent fragmentation
Fig. 1 depicts the time evolution of the orbital occupation. The figure shows that initially $(t<0)$ the system is fragmented for all choices of orbital number. The occupation in the natural orbitals changes quite rapidly after the quenching procedure. The occupation in the orbitals with initial significant population decreases, while the one in the initially nearly empty orbitals increases. This trend continues until a saturation point, located approximately at time $t=0.5$ for all values of $M$, where the population of the orbitals coalesces around the value $1/M$. This represents an irregular behavior which we term violent fragmentation, and is the hallmark of a complete saturation of the probed Hilbert space of our quantum simulations. We remark that this type of violent fragmentation is exhibited — practically unchanged — for all values of $M$ up to $M=20$, which represents a computational upper threshold for our simulations given the current state-of-the-art. At the transition time the fully collapsed state always occupies the full available configuration space. The quench from strongly repulsive to strongly attractive long-range interactions can therefore be described as a transition process from quantum to classical behavior of the long-range interacting bosons.
@@ 4.1 Violent fragmentation
4.2 Fock space delocalization
@@
To further understand the quantum to classical behavior, we visualize the value of the coefficients $C_{\bar{n}}(t)$ after the quench in Fig. 2 We plot $|C_{\bar{n}}(t)|^{2}$ as a function of the index $n$ of the basis states for the final simulation time at $t=3.0$. The number of configurations in the available space for particular choice of $M$ is determined by $N_{\mathrm{conf}}$ by Eq.(4). Thus, $N_{\mathrm{conf}}$ increases exponentially from 126 for $M=5$ to over 40,000 for $M=20$. For each choice of $M$, the initial TG gas undergoes a violent delocalization process in which most of the configurations acquire a comparable weight. This behavior corresponds to the definition of classical gas, which occupies the whole available configuration space.
@@
Figure 2: Configuration space population at time $t=3.0$ as measured from the amplitude squared of the many-body coefficients, for $N=5$ quenched bosons with $\alpha=3.0$, $g_{d}=100.0\to-100.0$ and increasing orbital number $M$. The orbital number is (a) $M=5$, (b) $M=10$, (c) $M=15$, (d) $M=20$.
4.3 Unbounded Entropy measures
@@
In the dynamical evolution of isolated quantum system, statistical relaxation is established by some kind of equilibrium. For interacting many-body systems, statistical relaxation is also related to chaos in the energy spectra [101]. In particular, chaos in quantum interacting systems is defined by the statistics of the eigenstates. The onset of chaos is established by quadratic growth in entropy at very short time, then linear growth before saturation to a maximum entropy states. For time reversal and rotationally invariant systems, the maximum entropy is also determined by the predictions of a Gaussian orthogonal ensemble of random matrices(GOE) [102, 47, 103, 104].
@@
Therefore, to further understand the violent dynamics and its impact on the relaxation process, in Fig 3 we plot the four different entropic quantities introduced in section 3 ($S_{C}(t)$, $S_{n}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$) as measures of irregular dynamics. We present how the different measures approach the GOE prediction as determined by the number of orbitals used in the computation. For a GOE of random matrices, $I_{C}^{GOE}=D/3$ and $S_{C}^{GOE}=\ln 0.48D$, where $D\times D$ is the dimension of the random matrices [102]. To compare with GOE estimates, we set $D$ as equal to the number of configurations $N_{\mathrm{conf}}$ participating in the dynamics. For occupation entropy $S_{n}$, $D$ is set equal to the number of orbitals $M$ used in the computations. Thus $S_{n}^{GOE}$ = $-\sum_{i=1}^{M}\frac{1}{M}\ln(\frac{1}{M})=\ln(M)$. The GOE estimates are evaluated for different $M$ used in the MCTDH-B treatment.
@@
Fig. 3 shows that statistical relaxation is achieved simultaneously with Fock space delocalization and violent dynamical fragmentation at time $t$=$0.5$. In particular, the results for increasing orbital number further support the thesis that the production of entropy is indeed unbounded. For example, for $M=15$, the number of configurations in the dynamics is $N_{conf}$ = 11’628. The corresponding GOE predictions are $S_{C}^{GOE}$ = 8.627, $S_{n}^{GOE}=2.708$, $I_{C}^{GOE}$= 3’876. We observe that for the prequench state $(t=0)$, all the measures are far enough from 00 indicating that the system initially is fragmented and delocalized. With time, all the entropy measures smoothly increase and then saturate. The saturation values for $M=15$ are $S_{C}^{\mathrm{sat}}\approx 8.94$, $S_{n}^{\mathrm{sat}}\approx 2.71$ and $I_{C}^{\mathrm{sat}}=5^{\prime}820$. These values are indeed close to GOE predictions, however a slight discrepancy exists because the interaction strength needed to exhibit a TG-like limit in the initial state is very large, albeit not infinite, which is the perquisite for GOE prediction. Repeating the same calculations with $M=20$ yields $N_{\mathrm{conf}}$=42’504, and the corresponding GOE predictions are $S_{C}^{GOE}$ = 9.923, $S_{n}^{GOE}=2.995$, $I_{C}^{GOE}$= 14’168. The corresponding saturation values in MCTDH-B computations are $S_{C}^{\mathrm{sat}}\approx 10.24$, $S_{n}^{\mathrm{sat}}\approx 2.99$ and $I_{C}^{\mathrm{sat}}\approx 21^{\prime}372$. The saturation close to GOE prediction for progressively larger values of $M$ (and hence the configuration space) indeed exhibits the hallmark of statistical relaxation. However, the impact of violent fragmentation is visible in the increase of the entropy measures with the number of orbitals. These combined results suggest that the entropy can increase without bound with increase of the orbital number, similarly to how a classical gas occupies the entire available configuration space when unconstrained.
@@
5 Understanding of relaxation process
@@
An intriguing question to understand is the intermediate path from a relaxing state to a fully relaxed one. In this section, we thus analyze the many-body quantum dynamics over the entire range of the exponents $\alpha\in\left[0.5,4\right]$. To understand the relaxation process we thoroughly discuss two typical cases: $\alpha=2.7$ (weakly long-range) and $\alpha=0.7$ (strongly long-range). Besides thoroughly investigating the dynamics of bosons, we also present calculations for spinless fermions to establish the universality of the fragmentation and relaxation processes. For both quantum statistics, we maintain the same protocol in the quench dynamics. We finally summarize the entire results in a parameter-time diagram which depicts the phase boundary between non-relaxed gas and relaxed gas.
5.1 Process of relaxation for bosons
@@
First, we discuss the case for strong long-range interactions with $\alpha=0.7$ and measure all the four entropic quantities $S_{C}(t)$, $S_{n}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$, which are plotted in Fig. 4. We observe a concurring estimate from all the four different measures that the relaxation happens around time $t\approx 0.6$. As expected, the saturation values approach the corresponding GOE predictions. Both $S_{c}(t)$ and $S_{C}^{N}(t)$ exhibit peculiar oscillations before reaching the saturation value. This behavior suggests a highly complex dynamics with several intermediate and rather unstable phases before the entropy eventually settles to the final state. $S_{n}$, instead, exhibits an almost linear increase at short times and smoothly settles to the GOE value. $S_{n}$ is more insensitive to irregularities in the dynamics as it is determined by the fragmentation of the many-body states, which is a many-body invariant. In fact, in the previous section we already saw that $S_{n}$ is the entropic quantity that best converges to the GOE result.
@@
The results for weak long-range interactions with $\alpha=2.7$ are presented in Fig. 5. The corresponding relaxation process is less complex than for strong long-range interactions. In the intermediate dynamics, the entropy measures exhibit a single but deeper minimum which is superseded by the eventual saturation to the GOE value. There is no clear signature of quasi equilibrium states for this case, and the time to reach the relaxed state is shorter ($t\approx 0.4$) than for strong long-range interactions.
@@
In Fig. 6, we summarize our results in the form of a parameter-time diagram across the interaction power $\alpha$ and the simulation time. Due to the closest agreement with GOE values for $S_{n}$ and its more regular dynamics that allows to determine the relaxation time more precisely, we will consider it as a good metric and use it as an order parameter in the parameter-time diagram. The diagram shows a clear separation between non-relaxed and relaxed states and we observe an exponential boundary between these two regimes. This suggests very long relaxation times in the limit $\alpha\to 0$, which cannot be reached with our current numerics due to the need of employing increasingly larger system sizes  [^3] We remark that the reduction of entropy exhibited by $\alpha=0.5$ at $t>0.6$ is likely due to self-interference of the long-range tails of the interacting particles. In our numerics, although we employ a harmonic confinement, the system is always defined with periodic boundary conditions. If the system size is not large enough, as it might have happened for the $\alpha=0.5$ results, self-interference patterns across the system boundaries will appear and modify the dynamics.. The parameter-time diagram has more distinct characteristics when $\alpha<1$, i.e. when the range is less than the dimension of the system. For larger values, instead, the exponential boundary gradually loses its structure and the relaxation time is almost independent of the choice of $\alpha$.
Figure 6: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched bosons with $g_{d}=100.0\to-100.0$ and $M=10$. The dashed line indicates the exponential boundary between evolving and fully relaxed states.
5.2 Process of relaxation for fermions
@@
Figure 9: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched fermions with $g_{d}=100.0\to-100.0$ and $M=10$. The dashed line indicates the exponential boundary between evolving and fully relaxed states.
6 Control of relaxation
@@
It is an established fact that a (near-)integrable system relaxes in at least two time scales related to the faster and slower relaxation processes [105]. Prethermalization occurs in the intermediate time scale before asymptotic relaxation to an equilibrium state. A necessary condition for thermalization is statistical relaxation in various observables to some kind of equilibrium. However, the observed dynamics in the sudden quench probed here is a highly nonintegrable phenomenon. The dynamics is so complex that different intermediate states may exist, but we can not detect them due to their extremely small lifetimes. It is thus impossible to establish the possibility of any prethermal state in the strongly quenched regime within the current computational limitations.
@@
Nevertheless, another key question that can be addressed is whether it is possible to control the complex and extremely nontrivial pathways observed in the relaxation dynamics during the quench by probing regimes of weaker interactions. In the following, we thus present an extended analysis of the dynamics for quenches with $g_{d}=+100.0$ to $g_{d}=-80.0,-50.0,-30.0$ and $-10.0$. Again, we evaluate the measures of entropy dynamics for $N=5$ bosons for two distinct choices of $\alpha$. We present the results for $\alpha=0.7$ in Fig. 10 and for $\alpha=2.7$ in Fig. 11. For $\alpha=0.7$, we observe that the dynamics is strongly affected by decreasing the interactions to weaker values. The system exhibits more complex pathways and the relaxation process is severely hampered. In the same time scale, when $g_{d}$ is quenched from $+100.0$ to $-100.0$ and the system exhibits perfect relaxation after passing through many complex intermediate states, the $g_{d}$= $+100.0$ to $-10.0$ quench is unable to achieve a completely relaxed state.
It is known that a linear increase in entropy is the precursor of thermalization [106]. In Fig. 10 (b), we indeed observe that $S_{n}(t)$ exhibits a linear increase when the interaction strength is quenched to the same magnitude. Gradually weakening the quench, though, the linear increase starts to disappear and instead an initial slower ramp starts to develop. For $g_{d}=+100.0$ to $g_{d}=-10.0$ quench, the growth of $S_{n}(t)$ clearly exhibits three different time scales rather than a linear increase of the entropy. In the first interval (I, green background), there is a clear quadratic increase which ends up in a very short plateau. In the second interval (II, yellow background), $S_{n}(t)$ increases more rapidly, acquiring the familiar linear trend of the stronger quenches. In the third interval (III, red background), the growth of $S_{n}(t)$ slows down and does not quite reach the GOE saturation value. Our findings are in perfect agreement with what observed in works dealing with two-body random interaction models described by superpositions of mean-field states that are used to describe the onset of chaotic dynamics [107]. In these models, the predicted entropy dynamics follows an initial quadratic growth, followed by a linear growth, before reaching eventual saturation. The quadratic entropy growth is the more generic case in systems where the spreading function follows a Breit-Wigner form (Lorentzian). This Lorentzian shape arises when the interaction strength is such that the spreading width $\Gamma$ is small compared to the energy range over which the states are spread. The linear regime, instead, appears when the interactions are strong enough to cause the spreading function (a.k.a. the local density of states) to take a Gaussian form because many more states become populated (as it is the case in Fig. 2). The linear growth of entropy in this regime reflects the system’s chaotic behavior, where the many-body state explores the energy space more uniformly.
The slower entropy increase and the different functional behavior than the immediate linear increase followed by saturation encountered in the strong quenches suggests that the entropy evolution can generally be manipulated by playing around with the strength and reach of the interactions. This finding is in agreement with the notion that the presence of long-range forces, such as in gravitational or Coulombic systems, tends to slow down the relaxation towards equilibrium due to the non-local nature of the interactions [108, 109, 110, 111, 13]. Thus, we conjecture that the entropy might be even slowed down entirely into a prethermal plateau following a similar interplay of interaction range and sudden quenches in related systems.
@@
Figure 10: Dynamics for four different entropy measures for $N=5$ quenched, strongly long-range interacting bosons with $\alpha=0.7$ and varying quench strength. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$. For all curves $M=10$ orbitals were used. In panel (b), the three different background colors denote the three different regimes observed in the $g_{d}=-100.0\to+10$ quench.
@@
The dynamics shown in Fig. 11 for shorter range interactions and smaller quench is distinctly different. The system remains mostly insensitive to the reduction in the quench strength. Some modulations in the amplitude are observed, but the basic conclusion remains unchanged: the system saturates to the GOE value in a single time scale, irrespective of the post-quench parameters. Fig 11(b) shows that $S_{n}(t)$ maintains its linearity at all times which guarantees thermal-like (i.e. linear) relaxation for all weaker quench processes. No appreciable dynamics slowdown is observed. Notably, the saturation value is almost independent of the postquench parameter. Even in this case, though, we can relate our findings to Ref. [107]. In the stronger quenches, there is a lot of energy injected into the system that make the particles interact very strongly from the get-go. As a result, only the linear regime and the eventual saturation can be detected. The quadratic regime should still exist, but it is pushed to very small times that do not appear very visible over longer time scales.
We thus conclude that when $\alpha$ is smaller than the dimension of the system, the dynamics can be drastically modified by the strength of the quench protocol, whereas for shorter range interactions the dynamics remains rather insensitive to it.
@@
7 Conclusions and Outlook
Understanding non-equilibrium dynamics and relaxation of quantum many-body systems is one of the major unsolved problems in quantum mechanics. The scarcity of appropriate theoretical tools to probe the relaxation dynamics makes the scenario even more challenging. When a generic isolated quantum system relaxes, the process of relaxation is not always known. There may be partial relaxation, or different stages of relaxation that can happen on different time scales leading to very complex nonequilibrium dynamics.